the fraction of defective items in a large lot is ‘p’. to test the null hypothesis h0: p=0.2, one consider the number ‘f’ of defectives in a sample of 8 items and accept the hypothesis if f≤6, and reject the hypothesis otherwise. what is the probability of type-1 error of the test? what is the probability of type-2 error of the test?




1
Expert's answer
2022-02-03T14:03:36-0500

The number of defectives "(f)" in the sample is a random variable which follows a Binomial distribution with parameters "n=8" and "p".

The probability of "x" defectives is given as,

"\binom{8}{x}p^x(1-p)^{8-x}"  


"a)"

The probability of type 1 error.

From definition, probability of type 1 error is he probability of rejecting "H_0" when it is true. So, "p=0.2" and we find the probability of 7 or 8 defectives.

So, when the number of defectives is  7, its probability is, "\binom{8}{7}0.2^7(0.8)^{1}=0.00008192"  and when the number of defectives is 8, its probability is, "\binom{8}{9}0.2^88^{8-8}=0.2^8=0.00000256"

The probability of committing type 1 error is 0.00008192+0.00000256=0.00008448


"b)"

The probability of type 2 error.

From definition, type 2 error is the probability of accepting "H_0" when "H_1" is true. So we determine the probability of 6 or less defectives when "p=0.1".

"p(x\le6)=1-p(x\gt 6)=1-\{\binom{8}{7}0.1^70.9+0.1^8\}=1-(7.3e-7)=0.99999927"Therefore, the probability of committing type 2 error is 0.99999927